OpenStudy (anonymous):

MEDAL . .. . Tell whether the given number is a solution of each equation: 87. 17=37+4f;f=-5

OpenStudy (sweetburger):

plug in -5 for f in the equation that will tell you

OpenStudy (sweetburger):

17 = 37 + 4(-5) are these statements equivalent?

OpenStudy (anonymous):

I really have no idea... I'm changing it to a medal AND a fan...

OpenStudy (nikkibxox):

It is. 17=37+4(-5) 17=37+(-20) 17=37-20 17=17 So yes. It is a solution